ABC manufactures a product with the following:

Direct materials standard 5 units at $4 per unit

Direct manufacturing labor standard 15 hours at $20 per hour

During the second quarter, the company made 3,000 products and used 22,000 units of raw material costing $90,200. Direct labor totaled 70,000 hours for $1,365,000.

Required:

a. Compute the direct materials price and efficiency variances for the quarter.

b. Compute the direct manufacturing labor price and efficiency variances for the quarter.
Answer #1

(a) Direct Material Price Variance = (Std Rate – Actual Rate) * Actual Quantity

Std Rate = $4 per unit

Actual rate = 90200/22000 = $ 4.1 per unit

Actual Qty = 22000 units

(4 – 4.1) * 22000 = 2200(U)

Direct Material Efficiency Variance = (Std Qty – Actual Qty) * Std Rate

Std Qty = 3000 * 5 = 15000 units

=(15000 – 22000) * 4 = 28000(U)

Direct Manufacturing Labour Price Variance = (Std Rate – Actual Rate) * Actual Hrs

Std Rate = $20 per hr

Actual rate = 1365000/70000 = $19.5 per hr

Actual Hrs = 70000 hrs

=(20 – 19.5) * 70000 = 35000(F)

Direct Manufacturing Labour Efficiency Variance = (Std Hrs – Actual Hrs) * Std Rate

Std Hrs = 3000 * 15 = 45000 hrs

(45000 – 70000) * 20 = 500000(U)